# Degree Distribution based Spiking Graph Networks for Domain Adaptation

### v2

@@ 4.2 Adversarial Distribution Alignment
$\mathcal{L}_{AD}=\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}\log Q\left(F(G_{i}^{s}),H(G_{i}^{s})|V_{th}^{D^{s}}\right)+\mathbb{E}_{G_{j}^{t}\in\mathcal{D}^{t}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}}\right)\right).$
@@
$\mathcal{L}_{AD}=$ $\mathbb{E}_{\begin{subarray}{c}G_{j}^{t}\in\mathcal{D}^{t}\\ D_{j}^{t}\subset{D^{s}}\end{subarray}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}}\right)\right)+\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}\log Q\left(F(G_{i}^{s}),H(G_{i}^{s})|V_{th}^{D^{s}}\right)$ (5)
@@
where $D^{t}=\{d_{i}^{t}|d_{i}^{t}\in D^{t},d_{i}^{t}\notin D^{s}\}$. Then, we iteratively update $V_{th}^{D^{t}}$ with Eq. 2 on each latency. Furthermore, we present an upper bound on the adversarial distribution alignment.
@@
$\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))\leq$ $\hat{\epsilon}_{S}(h,\hat{h}_{S}(\mathbf{S}))+2\mathbb{E}\left[\sup\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}$ (6)
$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right),$
@@
Theorem 2 proves the generalization bound of spiking graph domain adaptation. More details can be found in Appendix A.
@@ 4.3 Pseudo-label Distillation for Discrimination Learning
The goal of the pseudo-label distilling procedure is to keep those examples and their corresponding pseudo-labels from the deep feature space that aligns with the shallow feature space. Specifically, we denote $\mathbf{s^{\prime}}_{i}^{t}$ as the shallow spiking graph representation on the $L\textquoteright$ ’-th layer, where $L\textquoteright<L$, and $\hat{y}_{i}^{t}$ as the prediction of graph $G_{i}^{t}$ on the $L$-th layer.
Then, to enhance alignment between the shallow and deep feature spaces and facilitate the generation of more accurate pseudo-labels, we cluster the shallow features $\mathbf{s^{\prime}}^{t}$ into $C$ clusters and each cluster $\mathcal{E}_{j}$ includes graphs $\{G_{j}^{t}\}$.
After that, we find the dominating labels $e_{r}$ in the cluster, i.e., $\max_{r}|\{\mathcal{E}_{r}:e_{r}=\hat{y}_{j}^{t}\}|$, and remove other instances with the same pseudo-label but in different clusters. Formally, the pseudo-labels are signed with:
@@
where $H(\cdot)$ and $\mathbf{s}_{j}^{t}$ are the classifier and spiking graph representation, respectively, which are defined in Eq. 3. $l(\cdot)$ is the loss function, and we implement it with cross-entropy loss.
@@
$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)+\omega^{\prime},$
where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega^{\prime}=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is defined in Eq. 7.
The proof is detailed in Appendix B. From Theorem 3, we observe that the bound of DeSGDA is lower than simply aligning the distributions by incorporating the highly reliable pseudo-labels, demonstrating the effectiveness of pseudo labels for spiking graph domain adaptation.
@@ 4.4 Learning Framework
where $\lambda$ is a hyper-parameter to balance the adversarial training loss and classification loss. The learning procedure is illustrated in Algorithm C, and the complexity is shown in Appendix D .
@@
5 Experiment
5.1 Experimental Settings
Dataset. To demonstrate the effectiveness of DeSGDA, we conduct extensive experiments on four widely-used graph classification datasets from TUDataset  [^1] https://chrsmrrs.github.io/datasets/, including PROTEINS Dobson and Doig (2003), NCI1 Wale et al. (2008), FRANKENSTEIN Orsini et al. (2015), and MUTAGENICITY Kazius et al. (2005). To better address the variation in domain distributions within each dataset, we divided them into source and target domains based on the edge density, node density, and graph flux (i.e., the ratio of the number of nodes to the number of edges). The specific statistics, distribution visualization, and details introduction of experimental datasets are presented in Appendix E.
Baselines. We compare DeSGDA with competitive baselines on the aforementioned datasets, including one graph kernel method: WL subtree Shervashidze et al. (2011); four general graph neural networks: GCN Kipf and Welling (2017), GIN Xu et al. (2018), CIN Bodnar et al. (2021) and GMT Baek et al. (2021); two spiking graph neural networks: SpikeGCN Zhu et al. (2022) and DRSGNN Zhao et al. (2024); three recent domain adaptation methods: CDAN Long et al. (2018), ToAlign Wei et al. (2021b), and MetaAlign Wei et al. (2021a); and six graph domain adaptation methods: DEAL Yin et al. (2022), CoCo Yin et al. (2023), SGDA Qiao et al. (2023), DGDA Cai et al. (2024), A2GNN Liu et al. (2024b) and PA-BOTH Liu et al. (2024c). More details about the compared baselines can be found in Appendix F.
@@ 5.1 Experimental Settings
Implementation Details. DeSGDA and all baseline models are implemented using PyTorch [^2] https://pytorch.org/ and PyTorch Geometric [^3] https://www.pyg.org/. For DeSGDA, we deploy the GIN Xu et al. (2018) as the backbone of the degree-aware personalized spiking graph encoder, incorporating a mean-pooling layer for the readout function. We conduct experiments for DeSGDA and all baselines on NVIDIA A100 GPUs for a fair comparison, where the learning rate of Adam optimizer set to ${10^{-4}}$, hidden embedding dimension 256, weight decay ${10^{-12}}$, and GNN layers 4. Additionally, DeSGDA and all baseline models are trained using all labeled source samples and evaluated on unlabeled target samples Wu et al. (2020); Pang et al. (2023); Liu et al. (2023a); You et al. (2023). The performances of all models are measured and averaged on all samples for five runs.
[…]
SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
DeSGDA 68.5$\pm$1.2 71.4$\pm$1.3 70.1$\pm$0.7 69.0$\pm$1.1 68.9$\pm$1.0 66.3$\pm$1.4 69.6$\pm$1.3 70.2$\pm$1.7 71.1$\pm$1.6 69.3$\pm$1.8 74.4$\pm$1.6 70.0$\pm$1.9 70.1
[…]
DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
5.2 Performance Comparision
We present the results of the proposed DeSGDA with all baseline models under the setting of graph domain adaptation on different datasets in Table 1, 2, 3. From these tables, we observe that: (1) The performance of graph domain adaptation methods surpasses that of graph and spike-based graph methods. We attribute this improvement to the fact that domain distribution shifts degrade the performance of traditional graph methods. (2) The graph domain adaptation methods (DEAL and CoCo) outperform the spike-based graph methods (SpikeGCN and DRSGNN), underscoring the necessity of the research in spiking graph domain adaptation. (3) Our DeSGDA outperforms all baselines for most cases, which demonstrates its superiority over other methods. The remarkable performance of DeSGDA lies in two main reasons: (i) The degree-aware personalized spiking representations can capture more expressive information for graph classification by dynamically adjusting the thresholds of nodes in SNNs. (ii) The adversarial distribution alignment effectively addresses domain discrepancies by adversarially training the encoder and domain discriminator to align feature spaces. Moreover, the pseudo-label distillation aids in updating the source degree thresholds in the target domain, thereby ensuring optimal performance. More results evaluated on other datasets can be found in Appendix G.1.
Additionally, we conduct experiments to explore the flexibility of the proposed DeSGDA. Specifically, we replace the backbone of the degree-aware personalized spiking graph encoder (GIN) with various GNNs methods (i.e., GCN and GAT), and the results are shown in Figure 2. From the results, we observe that GIN consistently outperforms other GNNs architectures in most cases, demonstrating its powerful representation capability. This phenomenon also justifies our choice of using GIN to enhance the performance of the proposed DeSGDA. More results are reported in Appendix G.1.
5.3 Energy Efficiency Analysis
Figure 3: Energy consumption of DeSGDA on different datasets.
To assess the energy efficiency of DeSGDA, we use the metric from Zhu et al. (2022) and quantify the energy consumption for graph classification in the inference stage. Specifically, the graph domain adaption methods are evaluated on GPUs (NVIDIA A100), and the spiking-based methods are evaluated on neuromorphic chips (ROLLS Indiveri et al. (2015)) following Zhu et al. (2022). The results are shown in Figure 3, from the results, we find that compared with traditional graph domain adaptation methods, the spike-based methods (DeSGDA and DRSGNN) have significantly lower energy consumption, demonstrating the superior energy efficiency of SGNs. Moreover, although the energy consumption of DeSGDA is slightly higher than DRSGNN due to additional computations required for domain adaptation, the performance improvement justifies the deployment of DeSGDA in low-power devices.
@@ 5.3 Energy Efficiency Analysis
Table 4: The results of ablation studies on PROTEINS (source → target). Bold results indicate the best performance. Bold results indicate the best performance.
@@
DeSGDA w/ CL 74.9 73.3 73.7 75.1 77.0 71.3 73.9 70.0 78.1 77.6 78.8 74.9 75.1
DeSGDA 78.7 78.4 74.8 77.6 79.5 76.7 74.9 71.2 79.5 72.8 81.0 75.1 76.4
@@ 5.4 Ablation Study
We conduct ablation studies to examine the contributions of each component in the proposed DeSGDA: (1) DeSGDA w/o CA: It removes the adversarial distribution alignment module; (2) DeSGDA w/o PL: It removes the pseudo-label distilling module; (3) DeSGDA w/o CF: It removes the classification loss $\mathcal{L}_{S}$; (4) DeSGDA w/o TL: It utilizes the global thresholds on all nodes; (5) DeSGDA w/ PT: It deploys the adaptive perturbations Yin et al. (2022) on source data for alignment; (6) DeSGDA w/ CL: It replaces the adversarial learning with the cross-domain contrastive learning Yin et al. (2023).
Experimental results are shown in Table 4. From the table, we find that: (1) DeSGDA outperforms DeSGDA w/o CA, DeSGDA w/o PL, and DeSGDA w/o CF, demonstrating that the adversarial distribution alignment module can effectively reduce domain discrepancies, ensuring well-aligned feature spaces between source and target domains. Additionally, the pseudo-label distillation module can address the variance in thresholds across domains, while the classification loss $\mathcal{L}_{S}$ enables DeSGDA to effectively learn from labeled source data and generalize to the target domain. (2) DeSGDA w/o TL shows lower performance compared to DeSGDA, showing that the degree-aware thresholds, which are iteratively updated during model training, can resolve the issue of the inflexible architecture in SGNs. By using these thresholds, DeSGDA can effectively learn meaningful representations for nodes with various degrees. (3) DeSGDA outperforms DeSGDA w/ PT and DeSGDA w/ CL. We attribute that the adaptive perturbations method (DeSGDA w/ PT) can not effectively help DeSGDA overcome the domain discrepancy. Additionally, the cross-domain contrastive learning method (DeSGDA w/ CL) is less effective at aligning the source and target distribution compared to adversarial learning. More ablation results on other datasets are reported in Appendix G.2.

### v3

@@ 4.2 Adversarial Distribution Alignment
$\mathcal{L}_{AD}$ $=\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}\log Q\left(F(G_{i}^{s}),H(G_{i}^{s})|V_{th}^{D^{s}}\right)$
$+\mathbb{E}_{G_{j}^{t}\in\mathcal{D}^{t}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}}\right)\right).$
@@
$\mathcal{L}_{AD}$ $=\mathbb{E}_{\begin{subarray}{c}G_{j}^{t}\in\mathcal{D}^{t}\\ D_{j}^{t}\subset{D^{s}}\end{subarray}}\log\left(1-Q\left(F\left(G_{j}^{t}\right),H(G_{j}^{t})|V_{th}^{D^{s}}\right)\right)$
$+\mathbb{E}_{G_{i}^{s}\in\mathcal{D}^{s}}\log Q\left(F(G_{i}^{s}),H(G_{i}^{s})|V_{th}^{D^{s}}\right)$
@@
where $D^{t}=\{d_{i}^{t}|d_{i}^{t}\in D^{t},d_{i}^{t}\notin D^{s}\}$. We iteratively update $V_{th}^{D^{t}}$ with Eq. 2 on each latency. Furthermore, we present an upper bound for adversarial distribution alignment.
@@
$\epsilon_{T}(h,$ $\hat{h}_{T}(\mathbf{X}))\leq\hat{\epsilon}_{S}(h,\hat{h}_{S}(\mathbf{S}))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]$
$+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}+\min\Big{(}|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X}))|,$
$|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))|\Big{)}+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S},\mathbb{P}_{T}\right),$
@@
$\mathbf{P}_{kt}=\begin{array}[]{lr}\exp\left(\frac{u_{k}(t)-V_{th}}{\sigma(u_{k}(t)-u_{reset})}\right),&if\quad u_{\theta}\leq u(t)\leq V_{th},\\ 0,&if\quad u_{reset}\leq u_{k}(t)\leq u_{\theta}.\end{array}$ (5)
Theorem 2 proves the generalization bound of spiking graph domain adaptation. More details can be found in Appendix B.
@@ 4.3 Pseudo-label Distillation for Discrimination Learning
The goal of the pseudo-label distilling procedure is to keep those examples and their corresponding pseudo-labels from the deep feature space that aligns with the shallow feature space. Specifically, we denote $\mathbf{s^{\prime}}_{i}^{t}$ as the shallow spiking graph representation on the $L^{\prime}$-th layer, where $L^{\prime}<L$, and $\hat{y}_{i}^{t}$ as the prediction of graph $G_{i}^{t}$ on the $L$-th layer. Then, to enhance alignment between the shallow and deep feature spaces and facilitate the generation of more accurate pseudo-labels, we cluster the shallow features $\mathbf{s^{\prime}}^{t}$ into $C$ clusters and each cluster $\mathcal{E}_{j}$ includes graphs $\{G_{j}^{t}\}$. After that, we find the dominating labels $e_{r}$ in the cluster, i.e., $\max_{r}|\{\mathcal{E}_{r}:e_{r}=\hat{y}_{j}^{t}\}|$, and remove other instances with the same pseudo-label but in different clusters. Formally, the pseudo-labels are signed with:
@@
where $H(\cdot)$ and $\mathbf{s}_{j}^{t}$ are the classifier and spiking graph representation, which are defined in Eq. 3. $l(\cdot)$ is the loss function, and we implement it with cross-entropy loss.
@@
$\epsilon_{T}(h,\hat{h}_{T}($ $\mathbf{X}))\leq\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}\hat{\epsilon}_{T}(h,\hat{h}_{T}(S))+\frac{N_{S}}{N_{S}+N^{\prime}_{T}}\cdot\Bigg{(}\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S},\mathbb{P}_{T}\right)$
@@
$+\min\Big{(}|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X})))|,$
$|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X})))|\Big{)}\Bigg{)}\hat{\epsilon}_{S}(h,\hat{h}_{S}(S))+2\mathbb{E}\left[\sup\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\epsilon_{i}h(\mathbf{X}_{i},y_{i},p_{i})\right]$
$+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S},\mathbb{P}_{T}\right)+\omega^{\prime},$
where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega^{\prime}=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is defined in Theorem 2.
The proof is detailed in Appendix C. From Theorem 3, we observe that the bound of DeSGDA is lower than simply aligning the distributions by incorporating the highly reliable pseudo-labels, demonstrating the effectiveness of pseudo labels for spiking graph domain adaptation.
@@ 4.4 Learning Framework
where $\lambda$ is a hyper-parameter to balance the adversarial training loss and classification loss. The learning procedure is illustrated in Algorithm D, and the complexity is shown in Appendix E .
5 Experiment
5.1 Experimental Settings
Dataset. To demonstrate the effectiveness of DeSGDA, we conduct extensive experiments on four widely-used graph classification datasets from TUDataset  [^1] https://chrsmrrs.github.io/datasets/, including PROTEINS Dobson and Doig (2003), NCI1 Wale et al. (2008), FRANKENSTEIN Orsini et al. (2015), and MUTAGENICITY Kazius et al. (2005). To better address the variation in domain distributions within each dataset, we divided them into source and target domains based on the edge density, node density, and graph flux (i.e., the ratio of the number of nodes to the number of edges). The specific statistics, distribution visualization, and details introduction of experimental datasets are presented in Appendix F.
Baselines. We compare the proposed DeSGDA with competitive baselines on the aforementioned datasets, including one graph kernel method: WL subtree Shervashidze et al. (2011); four general graph neural networks: GCN Kipf and Welling (2017), GIN Xu et al. (2018), CIN Bodnar et al. (2021) and GMT Baek et al. (2021); two spiking graph neural networks: SpikeGCN Zhu et al. (2022) and DRSGNN Zhao et al. (2024); three domain adaptation methods: CDAN Long et al. (2018), ToAlign Wei et al. (2021b), and MetaAlign Wei et al. (2021a); and six graph domain adaptation methods: DEAL Yin et al. (2022b), CoCo Yin et al. (2023a), SGDA Qiao et al. (2023), DGDA Cai et al. (2024), A2GNN Liu et al. (2024a) and PA-BOTH Liu et al. (2024b). More detailed settings about the compared baselines are introduced in Appendix G.
@@ 5.1 Experimental Settings
Implementation Details. DeSGDA and all baseline models are implemented using PyTorch [^2] https://pytorch.org/ and PyTorch Geometric [^3] https://www.pyg.org/. For DeSGDA, we deploy the GIN Xu et al. (2018) as the backbone of the degree-aware personalized spiking graph encoder, incorporating a mean-pooling layer for the readout function. We conduct experiments for DeSGDA and all baselines on NVIDIA A100 GPUs for a fair comparison, where the learning rate of Adam optimizer set to ${10^{-4}}$, hidden embedding dimension 256, weight decay ${10^{-12}}$, and GNN layers 4. Additionally, DeSGDA and all baseline models are trained using all labeled source samples and evaluated on unlabeled target samples Wu et al. (2020). The performances of all models are measured and averaged on all samples for five different runs.
@@
Table 2: The results of ablation studies on PROTEINS (source → target). Bold results indicate the best performance.
@@
DeSGDA w/ CL 74.9 73.3 73.7 75.1 77.0 71.3 73.9 70.0 78.1 77.6 78.8 74.9 74.9
DeSGDA 78.7 78.4 74.8 77.6 79.5 76.7 74.9 71.2 79.5 72.8 81.0 75.1 76.7
5.2 Performance Comparision
We present the results of our proposed DeSGDA with all baseline models under the setting of graph domain adaptation on different datasets in Tables 1 and 13-23. From these tables, we observe that: (1) The performance of graph domain adaptation methods surpasses graph-based and spike-based graph methods. We attribute this improvement to the fact that domain distribution shifts degrade the performance of traditional graph methods. (2) The graph domain adaptation methods (DEAL and CoCo) outperform the spike-based graph methods (SpikeGCN and DRSGNN), underscoring the necessity of the research in spiking graph domain adaptation. (3) The WL subtree method outperforms SGDA, DGDA, A2GNN, and PA-BOTH but falls short compared to DEAL and CoCO. We attribute this to the relatively limited research specifically addressing the graph domain adaptation problem (e.g., DEAL and CoCo). To bridge this gap, we adapted node classification methods for graph classification tasks (e.g., SGDA, DGDA, A2GNN, and PA-BOTH). While the WL subtree method demonstrates superior performance over these adapted node classification-based methods, it remains less effective than dedicated graph domain adaptation methods tailored for graph classification tasks. (4) Our DeSGDA outperforms all baselines for most cases, which demonstrates its superiority over other methods. The remarkable performance of DeSGDA lies in two main reasons: (i) The degree-aware personalized spiking representations can capture more expressive information for graph classification by dynamically adjusting the thresholds of nodes in SNNs. (ii) The adversarial distribution alignment effectively addresses domain discrepancies by adversarially training the encoder and domain discriminator to align feature spaces. Moreover, the pseudo-label distillation aids in updating the source degree thresholds in the target domain, thereby ensuring optimal performance. More results on other datasets can be found in Appendix H.1.
Additionally, we conduct experiments to explore the flexibility of DeSGDA. Specifically, we replace the backbone of the degree-aware personalized spiking graph encoder (GIN) with various GNN methods (i.e., GCN and GAT), and the results are shown in Figure 2. From those results, we observe that GIN consistently outperforms other GNN architectures in most cases, demonstrating its powerful representation capability. This phenomenon also justifies our choice of using GIN to enhance the performance of the proposed DeSGDA. More results are reported in Appendix H.1.
5.3 Energy Efficiency Analysis
Figure 3: Energy consumption of DeSGDA and baselines on different datasets.
To assess the energy efficiency of DeSGDA, we use the metric from Zhu et al. (2022) and quantify the energy consumption for graph classification in the inference stage. Specifically, the graph domain adaption methods are evaluated on GPUs (NVIDIA A100), and the spiking-based methods are evaluated on neuromorphic chips (ROLLS Indiveri et al. (2015)) following Zhu et al. (2022). The results are shown in Figure 3, where we find that compared with traditional graph domain adaptation methods, the spike-based methods (DeSGDA and DRSGNN) have significantly lower energy consumption, demonstrating the superior energy efficiency of SGNs. Moreover, although the energy consumption of DeSGDA is slightly higher than DRSGNN due to additional computations required for domain adaptation, the performance improvement justifies the deployment of DeSGDA in low-power devices. Additionally, we present a comparison of training time and memory usage between DeSGDA and other graph domain adaptation methods. The results are detailed in Tables 9 and  10.
@@ 5.4 Ablation Study
We conduct ablation studies to examine the contributions of each component in the proposed DeSGDA: (1) DeSGDA w/o CA: It removes the adversarial distribution alignment module; (2) DeSGDA w/o PL: It removes the pseudo-label distilling module; (3) DeSGDA w/o CF: It removes the classification loss $\mathcal{L}_{S}$; (4) DeSGDA w/o TL: It utilizes the global thresholds on all nodes; (5) DeSGDA w/ PT: It deploys the adaptive perturbations Yin et al. (2022b) on source data for alignment; (6) DeSGDA w/ CL: It replaces the adversarial learning with the cross-domain contrastive learning Yin et al. (2023a).
Experimental results are shown in Table 2. From the table, we find that: (1) DeSGDA outperforms DeSGDA w/o CA, DeSGDA w/o PL, and DeSGDA w/o CF, demonstrating that the adversarial distribution alignment module effectively reduces domain discrepancies, ensuring well-aligned feature spaces between source and target domains. Additionally, the pseudo-label distillation module addresses variance in thresholds across domains, while the classification loss $\mathcal{L}_{S}$ enables DeSGDA to effectively learn from labeled source data and generalize to the target domain. (2) DeSGDA w/o TL shows lower performance compared to DeSGDA, showing that degree-aware thresholds, iteratively updated during training, resolve the issue of inflexible architecture in SGNs. By using these thresholds, DeSGDA effectively learns meaningful representations for nodes with various degrees. (3) DeSGDA outperforms DeSGDA w/ PT and DeSGDA w/ CL. We attribute this to the adaptive perturbations method (DeSGDA w/ PT) fail to help DeSGDA overcome domain discrepancies effectively. Additionally, the cross-domain contrastive learning method (DeSGDA w/ CL) is less effective at aligning the source and target distribution compared to adversarial learning. Furthermore, we provide ablation studies to examine the effect of directly replacing SGNs with commonly used Graph Neural Networks (GNNs) for generating representations for DeSGDA, and the results are shown in Tables 11, 12. More details about ablation results on other datasets are reported in Appendix H.3.